A particular first-order reaction has a rate constant of 1.35 × 102 s−1 at 25.0 ∘C. What is the magnitude of k at 75.0∘C if Ea = 55.5 kJ/mol. PLEASE SHOW WORK
Answer #1


Rate constant at 25.0°C, K, =1.35x10-51 Rate constant at 75.0°C, K2=? T, = 25.0°c =(273+ 25.0)K = 298K T =75.0°C =(75.0+ 273)